OpenStudy (anonymous):

An object with a massof 100 grams (0.100 kg) is dropped from a certain height and has a velocity of 60 meters/second. If the potential energy of this object is 179.99 joules, how high was the object when it was dropped?

OpenStudy (anonymous):

PE =m g h ---> h=PE/(mg)=179.99/(.1*9.81)= meter

OpenStudy (anonymous):

h=1892.7 m